Fruit drink A consists of 6% pure fruit juice and drink B consists of 15% pure fruit juice. How much of each kind of drink should you mix together to get 4 L of a 10% concentration of fruit juice?
Answer by mananth(16946)

percent percentage ---------------- quantity
Fruit juice I 6 ---------------- x liters
fruit juice II 15 ---------------- 4 - x liters
Mixture 10.00% ---------------- 4

6x+15(4-x)=10*4

6x+60-15x=40
6x-15x=40-60
-9x=-20
/-9
x= 2.22 liters 6.00% Fruit juice I
1.78 liters 15.00% fruit juice II